OpenStudy (anonymous):

Simplify expression: 5ln6-ln3

OpenStudy (anonymous):

Do you know how to do it without calculator

OpenStudy (anonymous):

5ln6-ln3=ln(6^5/3) 6ln2 +ln(x62-9)-ln(x+3)=ln((2^6*(62x-9))/(x+3… e^5ln3=ln(3^(e^5))

OpenStudy (anonymous):

yes i got it these is how to do without calculator it helps

OpenStudy (anonymous):

this how ru there @Lills

OpenStudy (lopus):